The electric field in a 1.6 mm x 1.6 mm square aluminum wire is    \(1.8 x 10^{-2} V/m\)   

What is the current in the wire?
Answer #1

R_of aluminum = 2.8*10^-8


Current density = 0.018 *10^8 / 2.8 = 6.428*10^5 A/m^2


Current = 1.6*1.6*10^-6 * 6.428*10^5 = 1.6455 A
Answer #2

current density J = i/A = E/rho

so i/A = E/rho

i = EA/rho

rho of Al    2.82 *10^-8

i = 0.018 * 1.6*1.6*10^-6/2.82*10^-8

i = 1.634 A